# A Decade of Human-Robot Interaction through Immersive Lenses: A Literature Review on Extended Reality as a Research Instrument in Social Robotics

### v1

@@ Abstract
Over the past decade, extended reality (XR), including virtual, augmented, and mixed reality, gained attention as a research instrument in human-robot interaction studies, but remains underexplored in empirical investigations of social robotics. To map the field, we systematically reviewed empirical studies from 2015 to 2025. Of 6,527 peer-reviewed articles, only 33 met strict inclusion criteria. We examined (1) how XR and virtual social robots are used and in which contexts, (2) data collection and analysis methods, (3) demographics of the researchers and participants, and (4) the stated challenges and future agendas. Our findings show that social XR-HRI research is still dominated by laboratory simulations, while crucial specifications like used hardware, software, and robots are often not reported. Robots typically act as passive and less interactive visual stimuli, while the rich biosignal (e.g., eye-tracking) and logging functions of modern head-mounted displays remain largely untapped. The research teams and samples are predominantly tech-centric, Western, young, and male, with frequent gaps in demographic reporting. Key limitations include hardware delays, small homogeneous samples, and short, shallow study cycles. We propose a five-phase roadmap to establish social XR-HRI as a reliable research medium, which includes fostering methodological innovation, a reinforced ecological validity by, e.g., using application contexts, the improvement of the robot’s interaction quality, promoting diversity in the sample and the development of a social XR-HRI taxonomy. Advancing in these directions is essential for XR to mature from a lab prototype into an ecologically valid research instrument for social robotics.
@@ 1 Introduction
Immersive technologies, including Virtual Reality (VR), Augmented Reality (AR), and Mixed Reality (MR) – collectively referred to as Extended Reality (XR) – have established themselves over the past decade as powerful research tools  [54]. In the field of human–robot interaction (HRI), XR has been applied across a wide spectrum of scenarios. These range from controlled studies examining social interactions between humans and robots  [62] to robot teleoperation  [41] and simulation settings where conducting experiments in the real world would be unsafe or impractical  [4]. Moreover, the flexibility and adaptability of XR technologies are increasingly leveraged to design and evaluate custom robot prototypes  [19]. Those virtual clones allow researchers to replicate a robot’s physical appearance and functional capabilities, while also enabling experimenting with prototypical designs or creative concepts that cannot yet be realized in physical robots due to technical or material limitations. The ability to rapidly modify robot design, behavior, and interaction modalities within virtual spaces shortens development cycles and enables studies that would be infeasible under real-world conditions  [9]. As a result, XR provides researchers with a highly dynamic simulation platform for exploring and studying robotic systems. Beyond its technical advantages, XR provides a means of examining the social dimensions of HRI in an environment that is both experimentally controlled and ecologically valid  [48]. Prior work has shown that human behavior in virtual environments often mirrors behavior in physical contexts  [23, 56], which increases confidence in the validity of findings from XR-based experiments  [32]. This makes XR particularly valuable in social robotics, where understanding human responses to robot gestures, speech, and other nonverbal cues is essential  [35]. Such interactions are not perceived solely on a functional level; rather, they activate human social schemas, prompting people to respond as if robots were social agents  [58, 57]. Prior studies demonstrate that even subtle variations of factors such as a robot’s anthropomorphic features, emotional expressiveness, or communication style can strongly shape user perceptions and trust [33]. Therefore a need for experimental flexibility and controllability regarding both the robot and its interaction context is required. Within the domain of social robotics, XR uniquely addresses these needs  [29], enabling researchers to conduct studies without access to physical robots, design entirely novel robotic embodiments, and test experimental communication channels - such as alternative speech patterns, gestures, or facial expressions. Moreover, contextual variables such as the study environment can be systematically manipulated to investigate their influence on human perception and behavior. Despite these opportunities and the growing number of XR-HRI studies, research that uses XR as a research tool to investigate social HRI remains fragmented across different application contexts. The variety of the conceptional roles of XR, the type of virtual robot embodiments, used measurements and technical gradations in terms of its quality makes is difficult to compare findings, assess external validity and to form a coherent picture of the field. To address this gap, this work presents a Systematic Literature Review (SLR) consolidating current knowledge, identifying recurring patterns, and outlining promising directions for future work of XR in social robotics. From this objective, the following research questions are derived:
@@
RQ1: What roles and contexts do XR and virtual robots occupy as a research tool for investigating social robots?
@@ 2 Background of XR in Research
XR has been established as a valid research instrument across various disciplines. For example, the medium is used in medicine as a training and simulation tool  [24], in educational settings to display content more immersive  [13], and as an awareness-raising tool for a wide range of social issues and inequalities  [28, 25]. Although experiments on the fundamental idea of virtual reality as we know it today were already conducted in the 1990s and early 2000s  [10, 5], XR only began to emerge as a growing factor in research from around 2015, following the release of advanced VR headsets from companies such as Facebook, HTC, and Sony. Since then, XR has evolved in terms of hardware, new application areas have been identified, and the advantages of XR as a research instrument have been recognized. These advantages include, for example, that virtual environments can present human simulations in a highly immersive way, creating a strong sense of presence. The concepts of presence and immersion in XR, identified by Rebelo et al. [55], have since been considered key principles, not only to be maintained in XR experiences but also as important factors in understanding the physical and psychological states of participants. Furthermore, presence plays a key role in the fact that VR, in particular, is widely acknowledged to offer high ecological validity. This means that researchers can have a highly controllable experimental environment  [12], while at the same time conducting studies in simulated field settings. It is also well established that people in VR display emotional responses similar to those in the real world  [23] and exhibit comparable behavior to real-world settings  [56]. All these advantages result in the possibility of conducting cost-effective, location-independent, and valid research across various fields.
@@ 3 Materials and Methods
This section reports the inclusion and exclusion criteria as well as the process of collecting relevant studies for this literature review following the guidelines proposed by  [37, 43, 47], as well as the coding scheme, which was used to systematically evaluates the relevant papers.
@@ 3.1 Inclusion and Exclusion Criteria
Studies were included if they addressed social HRI involving VR/AR/MR and explicitly employed these technologies as research instruments (e.g., simulation, experiment, or evaluation). Social HRI is defined as human–robot interaction that involves social communication or relationship-building processes (e.g., verbal/nonverbal cues, cooperation, trust, engagement), rather than purely functional task execution. Only empirical studies (e.g., experiments, user studies) published since January 2015 were considered. In addition, the articles had to be written in English and the full text of the article had to be accessible. Studies were excluded if they deviated from the topic of social robotics (e.g., industrial robots or no immersive medium) or if XR was only used for visualization without an experimental context. Furthermore, purely theoretical or technical articles without user evaluation were excluded, and care was taken to exclude outdated or non-generally available XR systems (e.g., experimental devices prior to 2015).
@@ 3.2 Identification of Articles
The literature review was conducted in IEEE Xplore, ACM Digital Library, Scopus, and Web of Science in April 2025, including articles published up to that date. The same search string was used in all databases, but adapted to the respective syntax: (("Human-Robot Interaction" OR "Robot" OR "HRI") AND ("Virtual Reality" OR "VR" OR "Mixed Reality" OR "MR" OR "Augmented Reality" OR "Extended Reality" OR "AR" OR "XR") AND ("Study" OR "Experiment" OR "Empirical" OR "Evaluation")). We ensured that the three core points – social-HRI, XR, and empirical studies – were covered. Using this search string, 6,527 peer-reviewed articles were found. Of these, 2,872 duplicate records were removed. The screening comprised two stages: In the first stage, titles and abstracts were checked according to the criteria described above, whereby 3,327 articles were not relevant; in the second stage, the full texts of the remaining contributions were reviewed. From 328 articles, 33 final articles remained, which entered the analysis process. The exluded articles were not primary research (e.g. technical papers), irrelevant topics (e.g. industrial robots) or not published in English. The PRISMA flow diagram, following the implications of  [47], is shown in Figure  1.
@@ 3.3 Coding Scheme
The unit of analysis was the individual paper. Coding was performed using a structured codebook, which contained operationalized definitions for all categories. For RQ1, codes were developed to classify the XR technology used, its overarching role, and the type of robot. XR is used in many settings, but it is rarely classified systematically. Cárdenas-Robledo et al. proposed an attempt at a taxonomy in the context of Industry 4.0 and note that XR roles often overlap and are difficult to distinguish clearly  [1]. They point out that multidimensional perspectives are needed to adequately describe XR applications. In the absence of alternative taxonomies, we adapted the categorization of Cárdenas-Robledo et al.  [1] and extended it with additional dimensions to describe robots (the type, the range of functions, whether Wizard-of-Oz (WoZ) was used, and the way of interaction). Accordingly, the codes shown in Table 1 were identified. All subcategories were created inductively, except those under Role/Contexts of XR, which were adopted from  [1]. For RQ2, codes for data collection and analysis methods were developed following Döring [14], and all additional codes belong to the research design were based on Creswell [11], as can be seen in Table 2. For RQ3, all subcategories under challenges and future work were created and categorized inductively (see Table 3). For RQ4, codes were developed deductively. Common demographic data such as discipline, status, gender, country, and age were collected for both researchers and participants. For researchers, this information was identified via search engines and may therefore not fully cover all aspects.
@@ 4.1 RQ1: Roles and Contexts of Extended Reality
The roles and contexts of XR in the reviewed studies are dominated by evaluation & testing and simulation. In about three-quarters of the papers, XR functions as a controllable laboratory environment for risk-free variation and measurement of robot concepts. Less common are roles such as training & education or prototyping & design, while robot–robot collaboration remains rare. XR is thus used mainly for observation rather than active co-creative interaction. Regarding the application domains, more than one third of studies examine fundamental HRI phenomena without a specific context. Among applied scenarios, education is most prominent, followed by healthcare, where XR supports safe, repeatable simulations for learning and rehabilitation. Public and commercial settings receive comparatively little attention. In terms of XR technologies, VR clearly dominates due to technical maturity and high experimental control. MR is emerging but remains limited by development complexity  [40]. AR plays only a minor role. Common devices include HTC Vive/Pro and Microsoft HoloLens 2, with Unity as the prevalent development framework. Some studies fail to report hardware/software details, which reduces reproducibility  [15]. Concerning the robot platforms, Pepper and NAO are most frequently used, while a few studies leverage XR to implement entirely virtual robots. Over one third rely on Wizard-of-Oz control, mainly to simulate naturalistic speech, meaning true autonomy is rare. Robot functionality is dominated by speech, gestures, and movement, typically through pre-defined utterances that ensure control but reduce ecological validity  [64]. Visual cues or tablet interactions are less common, and some robots serve merely as passive stimuli. With respect to interaction types, nearly half of the studies involve no participant–robot interaction at all. When interaction is allowed, it usually consists of short, pre-scripted speech or tablet commands. Only a small number of studies support dynamic speech interaction, mostly handled via Wizard-of-Oz mechanisms.
@@ 4.2 RQ2: Data & Analysis Collection and Research Design
Regarding data collection methods, questionnaires and observations are used most frequently, while interviews appear only occasionally. Physiological measurements remain rare despite the availability of integrated headset sensors. Real-time metrics such as eye tracking, pupillometry, or cognitive workload (e.g., via HP Reverb G2 Omnicept) are largely underutilized, and standardized psychological tests appear only in isolated cases. As a result, social XR-HRI research continues to rely mainly on subjective self-reports and manual behavioral coding. Concerning data analysis methods, all studies report descriptive statistics, and most employ inferential analyses. Comparative statistics appear in nearly half of the papers, while correlation and qualitative analyses are found primarily in mixed-methods designs. Applied research dominates, making up two-thirds of the studies; all works are empirical and analyze primary data. Most follow an explanatory approach, though exploratory components reflect the field’s novelty. Experiments are predominantly conducted in laboratory settings, which aligns with the XR focus. The few field studies identified all involve AR, suggesting AR is more suitable for real-world research than VR or MR. Regarding measurement timing, repeated-measures designs are most common, whereas cross-sectional and single-subject approaches occur less frequently. The limited use of longitudinal designs beyond basic repetitions indicates that temporal dynamics in social XR-HRI remain insufficiently explored.
@@ 4.3 RQ3: Current Challenges and Future Research
The most frequently reported challenges concern technical limitations and hardware. Restricted field of view, comfort issues, latency, and visual inconsistencies often disrupted immersion. Inaccurate position or depth tracking, network instability (e.g., UDP delays), and asynchronous timing between robot and display further reduced interaction quality. Software crashes, extensive calibration in WoZ setups, and the absence of haptic feedback or intuitive interfaces weakened usability. Overall, technology remains a major bottleneck, underscoring the need for robust hardware, stable networks, and XR-aware interaction design. Regarding methodological limitations, many studies questioned the ecological validity and real-world transferability of VR. Short tasks, narrow measurement windows, and reliance on subjective scales restricted behavioral insight. Sampling biases – driven by repeated-measure designs, technology-enthusiastic participants, and predominantly student populations – further limited generalizability. Constrained stimulus sets, limited modalities, and occasional multilingual data inconsistencies added complexity. These issues point to methodological fragility and highlight the need for mixed methods, proper control groups, and real-world validation. Sample size and representativeness posed additional challenges: many studies relied on small or pilot samples, with demographic imbalances (e.g., male-dominated, culturally homogeneous groups) and low XR experience producing novelty effects and limiting causal inference. Larger and more diverse samples are required to move beyond WEIRD (Western, Educated, Industrialized, Rich, Democratic) populations  [53]. Challenges also involved limited interaction depth and multimodality. Many studies relied on a single interaction channel despite participant demand for richer multimodal cues. Brief encounters prevented social bonding and reduced realism, while proxemics and gaze inconsistencies diminished authenticity. Technical constraints occasionally biased interaction logs, limiting ecological validity. Finally, studies often featured short durations and a lack of longitudinal data. Sessions lasting only minutes made it difficult to capture complex or slow-developing effects. Repetition in short intervals blurred distinctions between learning and technological impact, and the absence of long-term follow-ups left open questions about sustained behavioral change  [3].
@@
Future work aligns closely with these limitations, emphasizing technical advancements. Researchers call for better sensing and tracking (e.g., improved eye/hand tracking, biosignals), more stable networks, and lighter or HMD-free displays. Additional priorities include optimized GUIs, adaptive gesture generation, natural voice interaction, and support for complex cooperative tasks. Multi-user XR highlights the importance of synchronous interactions between multiple humans and robots. Deployment studies are recommended to align simulation and real-world outcomes. Testing in authentic environments (homes, schools, clinics) is considered essential for uncovering ecological factors such as lighting, noise, and connectivity that remain hidden in lab-based VR. Longitudinal studies are proposed to examine long-term learning, habituation, trust formation, and preference stability. Such designs are crucial for determining whether gains in motivation, competence, and trust endure over time and for deriving reliable design parameters for future robot systems. Social and emotional interaction is another key direction. Researchers seek to enhance robot nonverbal behaviors – gesture, emotion expression, gaze – while studying dialogue quality, trust-building, and relationship dynamics. Moving beyond task performance, the goal is to model authentic social bonding in XR. Improving measurements and metrics, including validated immersion scales, multidimensional pupil analyses, and longitudinal protocols, is considered essential for integrating technical, social, and methodological perspectives. Greater emphasis on diverse populations and cultural factors is needed to overcome the pronounced WEIRD bias and to understand culturally shaped interaction norms. Finally, adaptive and personalized systems and multimodal interaction are highlighted as mutually reinforcing directions. Real-time adaptation to user cognition, emotion, and context – potentially via reinforcement learning – depends on integrating multiple communication channels (speech, gestures, gaze). Together, they support richer and more accurate user modeling.
Table 3: Combined overview of codes and results from RQ3: Challenges and Future Work.

### v2

# A Decade of Human–Robot Interaction Through Immersive Lenses: Reviewing Extended Reality as a Research Instrument in Social Robotics

@@ Abstract
Over the past decade, Extended Reality (XR), including Virtual, Augmented, and Mixed Reality, gained attention as a research instrument in human-robot interaction studies, but remains underexplored in empirical investigations of social robotics. To map the field, we systematically reviewed empirical studies from 2015 to 2025. Of 6,527 peer-reviewed articles, only 33 met strict inclusion criteria. We examined (1) how XR and virtual social robots are used, focusing on the software and hardware employed and the application contexts in which they are deployed, (2) data collection and analysis methods, (3) demographics of the researchers and participants, and (4) the challenges and future directions. Our findings show that social XR-HRI research is still driven by laboratory simulations, while crucial specifications – such as the hardware, software, and robots used – are often not reported. Robots typically act as passive and hardly interactive visual stimulus, while the rich biosignal (e.g., eye-tracking) and logging (e.g. motion capturing) functions of modern head-mounted displays remain largely untapped. While there are gaps in demographic reporting, the research teams and samples are predominantly tech-centric, Western, young, and male. Key limitations include hardware delays, small homogeneous samples, and short study cycles. We propose a four-phase roadmap to establish social XR-HRI as a reliable research medium, which includes (1) strengthen application contexts, (2) more robust and testable technological iterations, (3) embedding diversity in samples and research teams, and (4) the need for reporting standards, e.g., in form of a suitable taxonomy. Advancing in these directions is essential for XR to mature from a lab prototype into an ecologically valid research instrument for social robotics.
@@ 1 Introduction
Immersive technologies, including Virtual Reality (VR), Augmented Reality (AR), and Mixed Reality (MR) – collectively referred to as Extended Reality (XR) – have been established over the past decade as powerful research tools [54]. In the field of human–robot interaction (HRI), XR has been applied across a wide spectrum of scenarios ranging from controlled studies examining social interactions between humans and robots [62] to robot teleoperation [41] and simulation settings where conducting experiments in the real world would be unsafe or impractical [4, 2]. Moreover, the flexibility and adaptability of XR technologies are increasingly leveraged to design and evaluate custom robot prototypes [19]. Those virtual clones allow researchers to replicate a robot’s physical appearance and functional capabilities, while also enabling experimenting with prototypical designs or creative concepts that cannot yet be realized in physical robots due to technical or material limitations. The ability to rapidly modify robot design, behavior, and interaction modalities within virtual spaces shortens development cycles and enables studies that would be infeasible under real-world conditions [9]. XR provides researchers with a highly dynamic simulation platform for exploring and studying robotic systems. Beyond its technical advantages, XR provides a means of examining the social dimensions of HRI in an environment that is both experimentally controlled and ecologically valid [48]. Prior work has shown that human behavior in virtual environments often mirrors behavior in physical contexts [23, 56], which increases confidence in the validity of findings from XR-based experiments [34]. This makes XR particularly valuable in social robotics, where understanding human responses to robot gestures, speech, and other nonverbal cues is essential [37]. Such interactions are not perceived solely on a functional level; rather, they activate human social schemas, prompting people to respond as if robots were social agents [58, 57]. Prior studies demonstrate that even subtle variations of factors such as a robot’s anthropomorphic features, emotional expressiveness, or communication style can strongly shape user perceptions and trust [35]. Therefore, a need for experimental flexibility and controllability regarding both the robot and its interaction context is required. Within the domain of social robotics, XR uniquely addresses these needs [30], enabling researchers to conduct studies without access to physical robots, design entirely novel robotic embodiments, and test experimental communication channels – such as alternative speech patterns, gestures, or facial expressions. Moreover, contextual variables such as the study environment can be systematically manipulated to investigate their influence on human perception and behavior. Despite these opportunities and the growing number of XR-HRI studies, research that uses XR as a research tool to investigate social HRI remains fragmented across different application contexts. The variety of (1) the conceptional roles of XR (e.g., simulation, prototyping, teleoperation, behavioral testing) and the interaction contexts in which social phenomena are investigated, (2) the methods and study design used to measure social cues, (3) the technical implementation and fidelity of virtual robots and environments, and (4) the composition of participant samples and research teams affects the comparability and generalizability of XR-HRI work. This heterogeneity makes it difficult to synthesize results, assess external validity, and derive a coherent picture of how XR is currently used to study social robots. To address this gap, this work presents a Systematic Literature Review (SLR) consolidating current knowledge, identifying recurring patterns, and outlining promising directions for future work on XR as a research tool in social robotics. From this objective, the following research questions are derived:
@@
RQ1: What roles and contexts do XR and virtual robots hold as a research tool for investigating social robots?
@@ 2 Background of XR in Research
XR has been established as a valid research instrument across various disciplines. For example, the medium is used in medicine as a training and simulation tool [25], in educational settings to display content more immersively [13], and as an awareness-raising tool for a wide range of social issues and inequalities [29, 26]. Although experiments on the fundamental idea of VR as we know it today were already conducted in the 1990s and early 2000s [10, 5], XR only began to emerge as a factor in research from around 2015, following the release of advanced VR headsets from companies such as Facebook, HTC, and Sony. Since then, XR has evolved in terms of hardware, new application areas have been identified, and the advantages of XR as a research instrument have been recognized. These advantages include, for example, that virtual environments can present human simulations in a highly immersive way, creating a strong sense of presence. The concepts of presence and immersion in XR, identified by Rebelo et al. [55], have since been considered key principles, not only to be maintained in XR experiences but also as important factors in understanding the physical and psychological states of participants. Furthermore, presence plays a key role in the fact that VR, in particular, is widely acknowledged to offer high ecological validity. This means that researchers can have a highly controllable experimental environment [12], while at the same time conducting studies in simulated field settings. It is also well established that people in VR display emotional responses similar to those in the real world [23] and exhibit comparable behavior to real-world settings [56]. All these advantages result in the possibility of conducting cost-effective, location-independent, and valid research across various fields.
@@ 3 Materials and Methods
This section reports the inclusion and exclusion criteria as well as the process of collecting relevant studies for this literature review following the PRISMA guidelines proposed by [43, 47], as well as the coding scheme, which was used to systematically evaluate the relevant publications.
@@ 3.1 Inclusion and Exclusion Criteria
Studies were included if they addressed social HRI involving VR/AR/MR and explicitly employed these technologies as research instruments (e.g., simulation, experiment, or evaluation). Social HRI is defined as human–robot interaction that involves social communication or relationship-building processes (e.g., verbal/nonverbal cues, cooperation, trust, engagement), rather than purely functional task execution. Only empirical studies (e.g., experiments, user studies) published since January 2015 were considered because since then, mass-market VR-devices have been introduced. In addition, the articles had to be written in English and the full text of the article had to be accessible. Studies were excluded if they deviated from the topic of social robotics (e.g., industrial robots or no immersive medium) or if XR was only used for visualization without an experimental context. Furthermore, purely theoretical or technical articles without user evaluation were excluded, and care was taken to exclude outdated or non-generally available XR systems (experimental devices prior to 2015, e.g., Oculus Rift DK1).
@@ 3.2 Identification of Articles
The literature review was conducted within the databases IEEE Xplore, ACM Digital Library, Scopus, and Web of Science in April 2025, including articles published up to that date. The same search string was used in all databases, but adapted to the respective syntax: (("Human-Robot Interaction" OR "Robot" OR "HRI") AND ("Virtual Reality" OR "VR" OR "Mixed Reality" OR "MR" OR "Augmented Reality" OR "Extended Reality" OR "AR" OR "XR") AND ("Study" OR "Experiment" OR "Empirical" OR "Evaluation")). We ensured that the three core points – social HRI, XR, and empirical studies – were covered. Using this search string, 6,527 peer-reviewed articles were found. Of these, 2,872 duplicate records were removed. The screening comprised two stages: In the first stage, titles and abstracts were checked according to the criteria described above, whereby 3,327 articles were not off-topic; in the second stage, the full texts of the remaining contributions were reviewed. Out of these 328 articles, 33 final articles remained, which entered the analysis process. The excluded articles were not primary research (e.g. technical papers), irrelevant topics (e.g. industrial robots) or not published in English. The PRISMA flow diagram, following the implications of [47], is shown in Figure 1.
@@ 3.3 Coding Scheme
The unit of analysis was the individual publication. Coding was performed using a structured codebook, which was created deductively and contained operationalized definitions for all categories. For RQ1, codes were developed to classify the XR technology used, its overarching role, and the type of robot. XR is used in many settings, but it is rarely classified systematically. Cárdenas-Robledo et al. proposed a taxonomy in the context of Industry 4.0, noting that XR roles often overlap and are difficult to distinguish clearly [1]. They point out that multidimensional perspectives are needed to adequately describe XR applications. In the absence of alternative taxonomies, we adapted the categorization of Cárdenas-Robledo et al. [1] and extended it with additional dimensions to describe robots (the type, the range of functions, whether Wizard-of-Oz (WoZ) was used, and the way of interaction). Accordingly, the codes shown in Table 1 were identified. All subcategories were created inductively, except those under role/context of XR, which were adopted from [1]. For RQ2, codes for data collection and analysis methods were developed following Döring [14], which provides established guidance for operationalizing and categorizing research methods. All additional codes belonging to the research design were based on Creswell [11], a widely used methodological guide to research design and its core design types, as can be seen in Table 2. For RQ3, all subcategories under challenges and future work were created and categorized inductively (see Table 3). For RQ4, codes were developed deductively. Common demographic data such as discipline, status, gender, country, and age were collected for both researchers and participants. For researchers, this information was identified via search engines.
@@ 4.1 RQ1: Roles and Contexts of Extended Reality
The roles and contexts of XR in the reviewed studies are dominated by evaluation & testing (n=27) and simulation (n=24). In about three-quarters of the papers, XR functions as a controllable laboratory environment for risk-free variation and measurement of robot concepts. Less common are roles such as training & education (n=8) or prototyping & design (n=8), while robot–robot collaboration (n=4) remains rare. XR is thus used mainly for observation rather than active co-creative interaction. Regarding the field of application, more than one third of studies examine fundamental HRI phenomena without a specific context (n=12). Among applied scenarios, education (n=7) is most prominent, followed by healthcare (n=7), where XR supports safe, repeatable simulations for learning and rehabilitation. Public and commercial settings (n=4) receive comparatively little attention. In terms of XR technologies, VR clearly dominates (n=20) due to technical maturity and high experimental control. MR is emerging (n=9) but remains limited by development complexity  [40]. AR plays only a minor role (n=5). Common devices include HTC Vive/Pro (n=7) and Microsoft HoloLens 2 (n=8), with Unity as the prevalent development framework (n=19), whereby some studies fail to report hardware/software details. Concerning the virtual robot platforms, Pepper (n=8) and NAO (n=6) are most frequently used, while a few studies leverage XR to implement entirely virtual robots (n=6). Over one third rely on Wizard-of-Oz (n=23) control, mainly to simulate naturalistic speech. Robot functionality is dominated by speech (n=20), gestures (n=15), and navigation (n=13), typically through pre-defined utterances that ensure control but reduce ecological validity  [48]. Visual cues or tablet (n=4) interactions are less common, and some robots serve merely as passive stimuli. With respect to interaction types, nearly half of the studies involve no participant–robot interaction (n=15) at all. When interaction is allowed, it usually consists of short, pre-scripted speech or tablet commands (n=10). Only a small number of studies support dynamic speech interaction, where participants could freely interact with the robot without having to adhere to predefined voice commands, mostly handled via Wizard-of-Oz mechanisms.
@@ 4.2 RQ2: Data & Analysis Collection and Research Design
Regarding data collection methods, questionnaires (n=29) and observations (n=25) are used most frequently, while interviews (n=10) appear only occasionally. Physiological measurements (n=7) remain rare despite the availability of integrated headset sensors. Real-time metrics such as eye tracking, pupillometry, or cognitive workload (e.g., via HP Reverb G2 Omnicept) are largely underutilized, and standardized psychological tests (e.g., Implicit Association Test) appear only in isolated cases (n=3). As a result, social XR-HRI research continues to rely mainly on subjective self-reports and manual behavioral coding. Concerning data analysis methods, all studies report descriptive statistics (n=33), and most employ inferential analyses (n=27). Comparative statistics (n=16) appear in nearly half of the papers, while correlation (n=10) and qualitative analyses (n=9) are found primarily in mixed-methods designs. Applied research dominates (n=22), making up two-thirds of the studies; all works are empirical (n=33) and analyze primary data (n=33). Most follow an explanatory (n=21) approach, though exploratory (n=15) components reflect the field’s novelty. Experiments are predominantly conducted in laboratory settings (n=29), which aligns with the XR focus. The few field studies (n=4) identified all involve AR. Regarding measurement timing, repeated-measures designs (n=14) are most common, whereas cross-sectional (n=10) and single-subject (n=9) approaches occur less frequently.
@@ 4.3 RQ3: Current Challenges and Future Research
The most frequently reported challenges concern technical limitations and hardware (n=21). Restricted field of view, comfort issues, latency, and visual inconsistencies often disrupted immersion. Inaccurate position or depth tracking, network instability (e.g., UDP delays), and asynchronous timing between robot and display further reduced interaction quality. Software crashes, extensive calibration in WoZ setups, and the absence of haptic feedback or intuitive interfaces weakened usability. Overall, technology remains a major bottleneck, underscoring the need for robust hardware, stable networks, and XR-aware interaction design. Regarding methodological limitations (n=21), many studies questioned the ecological validity and real-world transferability of VR. Short tasks, narrow measurement windows, and reliance on subjective scales restricted behavioral insight. Sampling biases – driven by repeated-measure designs, technology-enthusiastic participants, and predominantly student populations – further limited generalizability. Constrained stimulus sets, limited modalities, and occasional multilingual data inconsistencies added complexity. Sample size and representativeness (n=14) posed additional challenges: many studies relied on small or pilot samples, with demographic imbalances (e.g., male-dominated, culturally homogeneous groups) and low XR experience producing novelty effects and limiting causal inference. Challenges also involved limited interaction depth and multimodality (n=10), i.e., robotic systems that usually only have one modality of interaction and therefore appear to be less adaptive and dynamic. Many studies relied on a single interaction channel despite participant demand for richer multimodal cues. Brief encounters prevented social bonding and reduced realism, while proxemics and gaze inconsistencies diminished authenticity. Technical constraints occasionally biased interaction logs, limiting ecological validity. Finally, studies often featured short durations and a lack of longitudinal data (n=8).
@@
Future work aligns closely with these limitations, emphasizing technical advancements (n=19). Researchers call for better sensing and tracking (e.g., improved eye/hand tracking, biosignals), more stable networks, and lighter head-mounted displays. Additional priorities include optimized graphical user interfaces, adaptive gesture generation, natural voice interaction, and support for complex cooperative tasks. Multi-user XR highlights the importance of synchronous interactions between multiple humans and robots. Deployment studies (n=15) are recommended to align simulation and real-world outcomes. Testing in authentic environments (homes, schools, clinics) is considered essential for uncovering ecological factors such as lighting, noise, and connectivity that remain hidden in lab-based VR. Longitudinal studies (n=12) are proposed to examine long-term learning, habituation, trust formation, and preference stability. Such designs are crucial for determining whether gains in motivation, competence, and trust endure over time and for deriving reliable design parameters for future robot systems. Social and emotional interaction (n=11) is another key direction. Researchers seek to enhance robot nonverbal behaviors – gestures, emotion expression, gaze – while studying dialogue quality, trust-building, and relationship dynamics. Moving beyond task performance, the goal is to model authentic social bonding in XR. Improving measurements and metrics (n=8), including validated immersion scales, multidimensional pupil analyses, and longitudinal protocols, is considered essential for integrating technical, social, and methodological perspectives. Greater emphasis on diverse populations and cultural factors (n=7) was mentioned in order to overcome the pronounced WEIRD bias and to understand culturally shaped interaction norms. Finally, adaptive and personalized systems (n=5) and multimodal interaction (n=5) are highlighted as mutually reinforcing directions. Real-time adaptation to user cognition, emotion, and context – potentially via reinforcement learning – depends on integrating multiple communication channels (speech, gestures, gaze). Together, these measures support richer and more accurate user modeling.
Table 3: Overview of codes and results for RQ3: Challenges and Future Work.